OpenStudy (callisto):

Given \(f(x,y) = \sqrt[3]{xy}\), is \(\frac{\partial f}{\partial x}\) continuous at (0,0)?

OpenStudy (inkyvoyd):

wouldn't that be a yes? trying the limit for x constant and y constant gives the same result

OpenStudy (inkyvoyd):

I mean the limit for y taking x =0 is 0, and the limit for x taking y=0 is 0, because the cube root of a negative number remains negative

OpenStudy (callisto):

But it is \(\frac{\partial f}{\partial x}\), not f(x,y)?

OpenStudy (inkyvoyd):

*facepalm*

OpenStudy (inkyvoyd):

no because x^(2/3) becomes 0 and the expression approaches inf? inky is uselesss

OpenStudy (callisto):

Hmm.. Aren't you getting 0/0 though?

OpenStudy (inkyvoyd):

that indeterminant form implies discontinuity right?

OpenStudy (callisto):

But someone insists that it is continuous...

OpenStudy (inkyvoyd):

find da book's answer key?

OpenStudy (callisto):

No

OpenStudy (inkyvoyd):

when does l'hopital's rule apply?

OpenStudy (callisto):

infty / infty, or 0/0. But I just know the single variable case. Not sure of the multi-variable case.

OpenStudy (inkyvoyd):

wolframy alpha?

OpenStudy (ash2326):

Stumbles me as well, \[\lim_{x\to0, y\to 0} (\frac{y}{x^2})^\frac{1}{3}\]

OpenStudy (callisto):

I'm here to learn, not to ask for an answer from WolframAlpha.

OpenStudy (inkyvoyd):

therefore, I'm going to go figure out the right answer from Wolfram and see if they show steps

OpenStudy (callisto):

Only if you can figure that out.

OpenStudy (callisto):

@skullpatrol Mind giving a hand here?

OpenStudy (inkyvoyd):

it's not continuous, but no steps.

OpenStudy (asnaseer):

If you want to learn about this, then this might be useful: http://www.math.oregonstate.edu/home/programs/undergrad/CalculusQuestStudyGuides/vcalc/limcont/limcont.html

OpenStudy (asnaseer):

I basically states that: A function f of two variables is continuous at a point \((x_0,y_0)\) if: \(f(x_0,y_0)\) is defined \(\lim_{(x,y)\rightarrow (x_0,y_0)}f(x,y)\) exists \(\lim_{(x,y)\rightarrow (x_0,y_0)}f(x,y)=f(x_0,y_0)\)

OpenStudy (asnaseer):

*It basically...

OpenStudy (callisto):

Here's the problem. I let g(x,y) = \(\frac{\partial f}{\partial x}\) I thought when I needed to find the limit, it should be \(\lim_{(x,y)\rightarrow (0,0)}g(x,y)\) However, someone told me that it should be \(\lim_{x\rightarrow 0}(\lim_{y\rightarrow 0}g(x,y))\) I'm confused.

OpenStudy (asnaseer):

Look at the link I gave - it also has examples of how to calculate these limits

OpenStudy (callisto):

I know how to calculate the limits, I just don't understand what the right way to set up the limit is.

OpenStudy (asnaseer):

This video might help: http://www.youtube.com/watch?v=4-liutK41v0

OpenStudy (callisto):

Once again, the problem I have is to set up the limit, not to evaluate it. When we find \(\frac{\partial f}{\partial x}\), we assume y as a constant. Then we need to find the limit based on our assumption that y is a constant. In this case, is it still appropriate to set up the limit like \(\lim_{(x,y) \rightarrow (0,0)} \frac{\partial f}{\partial x}\) ?

OpenStudy (anonymous):

\[ \Large \frac{\partial f}{\partial x}=\frac{y}{3\sqrt[3]{x^2y^2}}=\frac{0}{0}\\ \Large \frac{\partial f}{\partial x}=\frac{y}{3\sqrt[3]{x^2y^2}}\frac{\sqrt[3]{x^4y^4}}{\sqrt[3]{x^4y^4}}=\frac{y\sqrt[3]{x^4y^4}}{3xy}=\frac{\sqrt[3]{x^4y^4}}{3x}=0 \]

OpenStudy (anonymous):

\[\Large \frac{\partial f}{\partial x}=\frac{y}{3\sqrt[3]{x^2y^2}}=\frac{0}{0}\\ \Large \frac{\partial f}{\partial x}=\frac{y}{3\sqrt[3]{x^2y^2}}\frac{\sqrt[3]{x^4y^4}}{\sqrt[3]{x^4y^4}}=\frac{y\sqrt[3]{x^4y^4}}{3xy}=\frac{\sqrt[3]{x^4y^4}}{3x}=\frac{\sqrt[3]{x}\sqrt[3]{y^4}}{3}=0 \]

OpenStudy (anonymous):

\[ \Large \frac{\partial f}{\partial x}=h(x,y)=\Large\frac{\sqrt[3]{x}\sqrt[3]{y^4}}{3} \]

OpenStudy (callisto):

\[\Large \frac{\partial f}{\partial x}=\frac{y}{3\sqrt[3]{x^2y^2}}\frac{\sqrt[3]{x^4y^4}}{\sqrt[3]{x^4y^4}}=\frac{y\sqrt[3]{x^4y^4}}{3x^2y^2}=...?\]

OpenStudy (callisto):

@cinar I think something's wrong after the second equal sign in the second line of your post.

OpenStudy (anonymous):

no, that's right..

OpenStudy (callisto):

\[\sqrt[3]{x^2y^2}\times \sqrt[3]{x^4y^4} = (x^2y^2)^\frac{1}{3}\times (x^4y^4)^\frac{1}{3} = [(x^2y^2)(x^4y^4)]^\frac{1}{3}=((xy)^6)^\frac{1}{3} = (xy)^2?\]

OpenStudy (anonymous):

you are right \[\Large \frac{\partial f}{\partial x}=h(x,y)=\Large\frac{\sqrt[3]{y}}{3\sqrt[3]{x^2}}\\ \Large x=rcost\\ \Large y=rsint\\ \Large \lim_{(x,y)\rightarrow (0,0)}h(x,y)=\frac{\sqrt[3]{y}}{3\sqrt[3]{x^2}}\\ \Large \lim_{r\rightarrow 0}\frac{\sqrt[3]{rsint}}{3\sqrt[3]{r^2cos^2t}}= \lim_{r\rightarrow 0}\frac{r^{\frac{1}{3}}\sqrt[3]{sint}}{3r^{\frac{2}{3}}\sqrt[3]{cos^2t}}=\lim_{r\rightarrow 0}\frac{r^{\frac{1}{3}}\sqrt[3]{sint}}{3r^{\frac{2}{3}}\sqrt[3]{cos^2t}}=\\ \Large \lim_{r\rightarrow 0}\frac{\sqrt[3]{sint}}{3r^{\frac{1}{3}}\sqrt[3]{cos^2t}}=0 \] not continuous at (x,y)=(0,0)

OpenStudy (anonymous):

\[ \Large \frac{\partial f}{\partial x}=h(x,y)=\Large\frac{\sqrt[3]{y}}{3\sqrt[3]{x^2}}\\ \Large x=rcost\\ \Large y=rsint\\ \Large \lim_{(x,y)\rightarrow (0,0)}h(x,y)=\frac{\sqrt[3]{y}}{3\sqrt[3]{x^2}}\\ \Large \lim_{r\rightarrow 0}\frac{\sqrt[3]{rsint}}{3\sqrt[3]{r^2cos^2t}}= \lim_{r\rightarrow 0}\frac{r^{\frac{1}{3}}\sqrt[3]{sint}}{3r^{\frac{2}{3}}\sqrt[3]{cos^2t}}= \Large \lim_{r\rightarrow 0}\frac{\sqrt[3]{sint}}{3r^{\frac{1}{3}}\sqrt[3]{cos^2t}}=\infty \]

OpenStudy (asnaseer):

@Callisto -- sorry I had to go eat something - am back now. Remember that for a limit to exist, it must converge to the same value no matter path you take towards that limit (assuming the path is within the domain of the function). Therefore, to calculate your limit:\[\lim_{(x,y)\rightarrow(0,0)}\frac{\sqrt[3]{y}}{3\sqrt[3]{x^2}}\]you could chose a path that lies along \(y=\alpha x^2\) where \(\alpha\) is some constant. This gives:\[\lim_{(x,y)\rightarrow(0,0)}\frac{\sqrt[3]{y}}{3\sqrt[3]{x^2}}=\lim_{(x,y)\rightarrow(0,0)}\frac{\sqrt[3]{\alpha x^2}}{3\sqrt[3]{x^2}}=\lim_{(x,y)\rightarrow(0,0)}\frac{\sqrt[3]{\alpha}}{3}=g(\alpha)\]i.e. the limit depends on the value of \(\alpha\). Therefore the limit does not exist.

OpenStudy (experimentx):

we are not approaching (0,0) from different directions. partial derivatives is a special case of directional derivative where y is constant or we only have a curve that is function of x. I think we should consider one particular value of alpha than general value of it.

OpenStudy (asnaseer):

Could use the definition of limit itself to get:\[\frac{\partial f}{\partial x}(0,0)=\lim_{h\rightarrow0}\frac{f(h,0)-f(0,0)}{h}\lim_{h\rightarrow0}\frac{0-0}{h}=\frac{0}{h}\]

OpenStudy (asnaseer):

*Could we use ...

OpenStudy (experimentx):

yes that what i believe to be correct.

OpenStudy (asnaseer):

Hmmm... So this implies that \(\large\frac{\partial f}{\partial x}\) IS continuous at (0,0)?

OpenStudy (experimentx):

i think it does ... there is a result on analysis that says that directional derivatives even if function is not continuous.

OpenStudy (asnaseer):

Interesting - looks like I need to do some reading :)

OpenStudy (experimentx):

i didn't solve it. @Callisto did it herself.

OpenStudy (asnaseer):

I awarded the medal on the basis that you gave me some insight into something I was unaware of :)

OpenStudy (callisto):

@experimentX is just being humble. He guided me and explained the ideas to me.

OpenStudy (asnaseer):

the main point here is we all learnt something :)

OpenStudy (experimentx):

lol ... still i learned something :)